How do I do this problem?
Eugene bought 27 comics and 12 magazines for $126. If 3 comics cost as much as 2 magazines, find the cost of each comic.

Answer by Maths68(1474)

Let
Cost of a Comic = c
Cost of a Manazine = m

Given
3comics = 2 Magazine
3c=2m
c=2m/3..........(1)
27 comics and 12 magazines for $126.
27c+12m=126...........(2)
Put the value of c from (1) to (2)
2m/3(27)+12m=126
2m*9+12m=126
18m+12m=126
30m=126
m=126/30
m=4.2
Put the value of m in (1)
c=2(4.2)/3
c=8.4/3
c=2.8

Cost of a Comic = c = $2.8
Cost of a Manazine = m = $4.2


Check
======
27c+12m=126
27(2.8)+12(4.2)=126
75.60+50.40=126
126=126